OpenStudy (anonymous):

hey! got a question with math, please help! Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)= x+a b g(x)=cx−d Part 2. Show your work to prove that the inverse of f(x) is g(x). Part 3. Show your work to evaluate g(f(x)). Part 4. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (mathmale):

f(x)= x+a g(x)=cx−d Your best bet is to go thru the actual steps of finding the inverse function of f(x), and then comparing the results to g(x). Steps: 1. Replace f(x) with "y" 2. Interchange x and y Please do these two things now; then we'll continue

OpenStudy (anonymous):

hey sorry, i logged of for a minute, okay.

OpenStudy (anonymous):

So y = x + a, and interchange? sorry,

OpenStudy (anonymous):

oh okay, for g(x)=cx−d, it would be x = cy - d?

OpenStudy (mathmale):

Let's concentrate of f(x) first. In y = x + a, where you see x, substitute y; where you see y , substitute x.

OpenStudy (anonymous):

oh, okay, so x = y + a?

OpenStudy (mathmale):

Yes, very good. Now, please solve x = y + a for y.

OpenStudy (mathmale):

(just subtract "a" from both sides of x = y + a)

OpenStudy (anonymous):

okay, minus X from both sides and you get Y = x - a

OpenStudy (anonymous):

oh!!!! okay,, i though you did x

OpenStudy (anonymous):

so x - a = y?

OpenStudy (mathmale):

yes, or y = x - a. Now, replace y with the symbol for "inverse funtion of f(x)." It is:

OpenStudy (mathmale):

\[f ^{-1}(x)\]

OpenStudy (anonymous):

okay, so f−1(x) = x - a

OpenStudy (anonymous):

oops, f^-1(x) = x - a

OpenStudy (mathmale):